1. A laptop manufacturer produces β€˜π‘₯’ sets per week at total cost of


𝑅𝑠. (2π‘₯/25+ 3π‘₯ + 100). He is a monopolist and the demand for his product


is: π‘₯ = 75 βˆ’ 3𝑝, where 𝑝 is the price in rupees per set. Shoe that the


maximum revenue is obtained when about 30 sets are produced per week.


What is the monopoly price?

1
Expert's answer
2022-02-20T16:02:06-0500

The maximum revenue TR is obtained when "TR'(x) = MR = (p\u00d7x)' = 25 - 2x\/3 = 0."

The monopoly price and quantity are:

MR = MC,

25 - 2x/3 = 2x/25 + 3,

56/75 x = 22,

x = 29.46 units,

p = 25 - 29.46/3 = 15.18.